# Coupled Twiss Parameters Estimation from TbT Data

### v2

@@ VI One-turn Matrix Reconstruction
Coupled Twiss parameters can be estimated from the one-turn matrix at selected BPM or its power $M^{k}=NR^{k}N^{-1}$. Since the coordinates and momenta at different turns at a BPM are related by
${(q_{x},p_{x},q_{y},p_{y})}_{n+k}=M^{k}{(q_{x},p_{x},q_{y},p_{y})}_{n}$ (26)
@@
The estimated matrix $\hat{M}^{k}$ is generally not symplectic. Therefore, symplectification is performed before calculating the Twiss parameters [18]
@@
where $S$ is the standard symplectic form. The matrix $\bar{M}^{k}$ serves as a symplectic approximation of the matrix $M^{k}$. This matrix can then used to determine the coupled Twiss parameters [19, 30].
@@ VII Complex Normalized Spectrum
Using the normalization matrix $N$ and coordinates $(q_{x},p_{x},q_{y},p_{y})$, one can calculate normalized coordinates as follows
$(Q_{x},P_{x},Q_{y},P_{y})=N^{-1}(q_{x},p_{x},q_{y},p_{y})$ (28)
In this normalized space, phase space trajectories form circles when correct normalization matrix $N$ is applied. Introducing the following complex coordinates
@@
Figure 2: The spectrum of oscillations for complex coordinates $W_{x}$ (in red) and $W_{y}$ (in blue) with mismatched (a.) and matched (b.) normalization matrix $N$. For the matched case, only the main amplitudes $a_{x,x}$ and $a_{y,y}$ are nonzero.
Transverse TbT signals $x$ and $y$ are available at each BPM. The frequencies of oscillations $\nu_{x}$ and $\nu_{y}$ can be estimated with high accuracy from these signals. In the linear approximation, if $N$ doesn’t perform full normalization, the spectrum of complex coordinates contains two peaks at frequencies $(\nu_{x},\nu_{y})$ and two peaks at reflected frequencies $(\bar{\nu}_{x},\bar{\nu}_{y})=1-(\nu_{x},\nu_{y})$. Figure 2 illustrates the use of both mismatched and matched normalization matrices in computing normalized spectra. In addition to the primary peaks at the main frequencies, additional peaks appear due to coupled oscillations for the mismatched case.
In the case of nonlinear motion, harmonics of the main frequencies will also be present, but for measurements of linear optical model parameters, nonlinear effects can usually be neglected. Given two main frequencies and two reflected frequencies, four amplitudes can be computed from each spectrum. For example, the amplitudes of peaks for horizontal normalized coordinates $W_{x}$ can be estimated as
@@
The optimization objective below can be minimized to determine the elements of the normalization matrix $N$
$\min_{N}{\left(\frac{b_{x,x}+a_{x,y}+b_{x,y}}{a_{x,x}}+\frac{a_{y,x}+b_{y,x}+b_{y,y}}{a_{y,y}}\right)}$ (31)
@@ VIII Coupled Linear Invariants
For linear motion, two invariants, $I_{x}$ and $I_{y}$, exist, also referred to as actions. To determine the coupled normalization matrix, the objective function below can be minimized at each BPM
$\min_{N,I_{x},I_{y}}\left({\left(Q_{x}^{2}+P_{x}^{2}-2I_{x}\right)}^{2}{\left(Q_{y}^{2}+P_{y}^{2}-2I_{y}\right)}^{2}\right)$ (32)
@@
Figure 3: Fitted invariant values for the horizontal (a.) and vertical (b.) planes. 512 random samples are used for each BPM. The true values of invariants are shown with black dashed line, the blue lines show estimated invariants values with corresponding uncertainties.
For each monitor, the values and variance of the invariants can be estimated, and a weighted average can be computed to estimate invariants over all BPMs. This allows to fix the invariant values. Another option is to use invariant values estimated by using measured amplitudes and either modeled or measured beta functions. Thus, to fit the elements of the matrix $N$ using fixed invariants $\hat{I}_{x}$ and $\hat{I}y$, the following objective function needs to be minimized
$\min_{N}\left({\left(Q_{x}^{2}+P_{x}^{2}-2\hat{I}_{x}\right)}^{2}{\left(Q_{y}^{2}+P_{y}^{2}-2\hat{I}_{y}\right)}^{2}\right)$ (33)
@@ IX Experimental Results at the VEPP-4M
Figure 4: Example of TbT signals for horizontal (a.) and vertical (b.) planes.
@@ IX.1 Uncoupled Twiss Parameters
In Figure 5, the measured and model phase advances between adjacent BPMs are shown. The measured phase advances were computed from TbT data using $512$ turns. Amplitudes were computed using the first 128 turns, where the effect of decoherence can be neglected (Figure 4). Using the amplitudes and measured beta functions, estimates of linear invariants were obtained at each BPM, and the weighted averages were used to fix the invariant values for the method described in Section VIII. Figure 6 shows the invariant values at each BPM and the corresponding weighted average values.
Figure 5: Comparison of the measured (in red) and model (in black) phase advances between adjacent BPMs for the horizontal (a.) and vertical (b.) planes.
Figure 6: Estimated horizontal (a.) and vertical (b.) linear invariants for each BPM (in red) and corresponding weighted averages with errors (dashed).
Figure 7: Comparison of model (in black), from amplitude (in red) and from phase (in blue) beta functions for horizontal (a.) and vertical (b.) planes.
Uncoupled Twiss parameters were computed using the amplitudes and phases of TbT signals. An average over several combinations, with outlier cleaning, is used to compute Twiss parameters from phase data. The comparison with model values is shown in Figure 7. On average, the discrepancy between measured and model values is around 10-20%. Values obtained by different methods also differ by around 15%. Figure 9 shows an example of phase space trajectory normalization (first 128 turns) using model (black) and measured from phase data (red) Twiss parameters. While the normalized horizontal trajectory is close to a circle for both modeled and measured Twiss parameters, the vertical trajectory is not properly normalized. To perform full normalization, coupling should be included.
@@ IX.2 Coupled Twiss Parameters
Figure 8 shows the estimated in-plane beta functions obtained using methods described in Sections VI, VII, and VIII. For comparison, beta functions estimated using the method based on TbT phases are also shown. All methods for estimating coupled Twiss parameters use reconstructed momentum, which was computed from LS fit using two monitors around the probed monitor and modeled transport matrices. In all cases, the first 128 turns were used from a single TbT measurement. The method based on the reconstruction of the one turn matrix differs most from the other methods. The method based on the spectrum of complex normalized coordinates and the method based on linear coupled invariants are in good agreement with the results obtained from phase measurements. This agreement also indicates that the Twiss parameters specific to coupling, estimated using these methods, are expected to be meaningful and can be used for coupling characterization.
@@
Figure 8: Comparison of in-plane beta function estimated using different methods based on momentum reconstruction for horizontal (a.) and vertical (b.) planes.
Using the full set of coupled Twiss parameters, one can build a coupled normalization matrix and use it to normalize phase space trajectories. Figure 9 shows an example of such normalization using both uncoupled and coupled normalization matrices. In both cases, the shape of the horizontal phase space trajectory is close to a circle after normalization. For the vertical phase space trajectory, in the case of the uncoupled normalization matrix, there is a strong deviation from a circle due to coupling. With the coupled normalization matrix, the shape of the vertical phase space trajectory is close to a circle. The latter is further evidence that the considered methods produce sensible results.
@@
Figure 9: Normalized phase space trajectory using the model normalization matrix (in black) compared with the trajectory using the measured normalization matrix (in red) for the uncoupled (a. and b.) and coupled (c. and d.) cases.
Figure 10: Comparison of coupled amplitudes for horizontal (a.) and vertical (b.) planes, computed directly from TbT data and using estimated linear invariants and coupled Twiss parameters.
@@
Another option to check the validity of the estimated coupled Twiss parameters is to compute the coupled amplitudes $a_{x,y}$ and $a_{y,x}$. These amplitudes can be computed directly from TbT data and using estimated linear invariants and coupled Twiss parameters as follows
@@
$a_{y,x}^{2}=2\hat{I}_{x}\left(n_{3,1}^{2}+n_{1,4}^{2}\frac{n_{3,3}^{2}}{n_{1,1}^{2}}\right)$ (34)
where $n_{i,j}$ are the elements of the normalization matrix defined in Section IV. The result of such a comparison is shown in Figure 10. A qualitative argument for the validity of all methods can be observed, which further indicates the correctness of the estimated coupled Twiss parameters. The results are more accurate for the vertical plane, since $a_{y,x}$ amplitudes are several times larger.
@@ X Summary
In this paper, we examine the theoretical foundations of several methods for estimating the Twiss parameters, taking into account the coupling of transverse oscillations. Methods for the estimation of coupled Twiss parameters can be accessed in the reference [37]. All the described methods can utilize either one or several TbT measurements. In all methods, the reconstructed momenta, which can be computed using several BPMs and either modeled or measured transport matrices between them, is used. Since the methods directly use TbT coordinates, they are sensitive to BPM calibration errors.
The method based on reconstructing the one-turn matrix or its power uses TbT coordinates and reconstructed momenta to fit the corresponding matrix. Symplectification procedure is performed on the estimated matrix prior to the computation of coupled Twiss parameters. Another method uses normalized complex spectra and a special form of the normalization matrix to minimize an objective that depends on coupling amplitudes. This minimization procedure provides an estimation of the coupled normalization matrix. In the last method, the objective function to be minimized involves linear coupled invariants. These invariants can be estimated from the minimization procedure, fixed by prior optimization, or computed by other means. Numerical simulations show high accuracy in determining the coupled Twiss parameters using the described methods, taking into account random BPM errors and model errors.
@@ XI Acknowledgments
This work was partially supported by the Ministry of Science and Higher Education of the Russian Federation within the governmental order for Boreskov institute of Catalysis (project FWUR-2022-0001).

### v3

@@ VI One-turn Matrix Reconstruction
Coupled Twiss parameters can be estimated from the one-turn matrix at selected BPM or its power $M^{k}=NR^{k}N^{-1}$. Since the coordinates and momenta at different turns at a BPM are related by:
${(q_{x},p_{x},q_{y},p_{y})}^{T}_{n+k}=M^{k}{(q_{x},p_{x},q_{y},p_{y})}^{T}_{n}$ (26)
@@
The estimated matrix $\hat{M}^{k}$ is generally not symplectic. Therefore, symplectification [23] is performed before calculating the Twiss parameters:
@@
where $S$ is the standard symplectic form. The matrix $\bar{M}^{k}$ serves as a symplectic approximation of the matrix $M^{k}$. This matrix can then be used to determine the coupled Twiss parameters [16, 35].
@@ VII Complex Normalized Spectrum
Using the normalization matrix $N$ and coordinates $(q_{x},p_{x},q_{y},p_{y})$, one can calculate normalized coordinates as follows:
$(Q_{x},P_{x},Q_{y},P_{y})^{T}=N^{-1}(q_{x},p_{x},q_{y},p_{y})^{T}$ (28)
In this normalized space, phase space trajectories form circles when correct normalization matrix $N$ is applied. Introducing the following complex coordinates:
@@
Figure 2: (Color) The spectrum of oscillations for complex coordinates $W_{x}$ (in red) and $W_{y}$ (in blue) with mismatched (a.) and matched (b.) normalization matrix $N$. For the matched case, only the main amplitudes $a_{x,x}$ and $a_{y,y}$ are nonzero.
Transverse TbT signals $x$ and $y$ are available at each BPM. The frequencies of oscillations $\nu_{x}$ and $\nu_{y}$ can be estimated with high accuracy from these signals [40, 41, 42]. In the linear approximation, if $N$ does not perform full normalization, the spectrum of complex coordinates contains two peaks at frequencies $(\nu_{x},\nu_{y})$ and two peaks at reflected frequencies $(\bar{\nu}_{x},\bar{\nu}_{y})=1-(\nu_{x},\nu_{y})$. Figure 2 illustrates the use of both mismatched and matched normalization matrices in computing normalized spectra. In addition to the primary peaks at the main frequencies, additional peaks appear due to coupled oscillations for the mismatched case.
In the case of nonlinear motion, harmonics of the main frequencies will also be present, but for measurements of linear optical model parameters, nonlinear effects can usually be neglected. Given two main frequencies and two reflected frequencies, four amplitudes can be computed from each spectrum. For example, the amplitudes of peaks for horizontal normalized coordinates $W_{x}$ can be estimated as:
@@
The optimization objective below can be minimized to determine the elements of the normalization matrix $N$:
$\operatorname*{arg\,min}_{N}{\left(\frac{b_{x,x}+a_{x,y}+b_{x,y}}{a_{x,x}}+\frac{a_{y,x}+b_{y,x}+b_{y,y}}{a_{y,y}}\right)}$ (31)
@@ VIII Coupled Linear Invariants
For linear motion, two invariants, $I_{x}$ and $I_{y}$, exist, also referred to as actions. To determine the coupled normalization matrix, the objective function below can be minimized at each BPM:
$\operatorname*{arg\,min}_{N,I_{x},I_{y}}\sum_{i}{\left(Q_{x,i}^{2}+P_{x,i}^{2}-2I_{x}\right)}^{2}+{\left(Q_{y,i}^{2}+P_{y,i}^{2}-2I_{y}\right)}^{2}$ (32)
@@
Figure 3: (Color) Fitted invariant values for the horizontal (a.) and vertical (b.) planes. 512 random samples are used for each BPM. Samples for each BPM are grouped by vertical lines. The true values of invariants are shown with black dashed line, the blue lines show estimated invariants values with corresponding uncertainties.
For each monitor, the values and variance of the invariants can be estimated, and a weighted average can be computed to estimate invariants over all BPMs. This allows to fix the invariant values. Another option is to use invariant values estimated by using measured amplitudes and either modeled or measured beta functions. Thus, to fit the elements of the matrix $N$ using fixed invariants $\hat{I}_{x}$ and $\hat{I}y$, the following objective function needs to be minimized:
$\operatorname*{arg\,min}_{N}\sum_{i}{\left(Q_{x,i}^{2}+P_{x,i}^{2}-2\hat{I}_{x}\right)}^{2}+{\left(Q_{y,i}^{2}+P_{y,i}^{2}-2\hat{I}_{y}\right)}^{2}$ (33)
@@ IX Experimental Results at the VEPP-4M
Figure 4: (Color) The VEPP-4M model beta functions and BPM locations. The beta function values at BPMs are denoted with black crosses.
@@
Figure 5: (Color) Example of measured TbT signals for horizontal (a.) and vertical (b.) planes obtained at STP4 monitor.
@@ IX.1 Uncoupled Twiss Parameters
In Figure 6, the measured and model phase advances between adjacent BPMs are shown. The measured phase advances were computed from TbT data using 512 turns. Amplitudes were computed using the first 128 turns, where the effect of decoherence can be neglected (Figure 5). Using the amplitudes and measured beta functions, estimates of linear invariants were obtained at each BPM. Figure 7 shows the invariant values at each BPM and the corresponding weighted average values (in red).
Figure 6: (Color) Comparison of the measured phase advance between adjacent BPM pairs with corresponding model values $(\Delta\mu^{\textrm{m}}-\Delta\mu)/\Delta\mu^{\textrm{m}}$ for the horizontal (in red) and vertical (in blue) planes.
Figure 7: (Color) Estimated horizontal (a.) and vertical (b.) linear invariants for each BPM from TbT amplitudes (in red), invariant fit (in blue), complex normalized spectrum (in black), and reconstructed $n$–turn matrix (in magenta). Corresponding weighted averages with errors are shown with dashed lines for values estimated from amplitude and invariant fit.
Figure 8: (Color) Comparison of the measured beta functions estimated from amplitude (in red) and phase (in blue) with corresponding model values $(\beta^{\textrm{m}}-\beta)/\beta^{\textrm{m}}$ for horizontal (a.) and vertical (b.) planes.
Uncoupled Twiss parameters were computed using the amplitudes and phases of TbT signals. An average over several combinations, with outlier cleaning, is used to compute Twiss parameters from phase data [30]. The comparison with model values is shown in Figure 8. The discrepancy between measured and model values is around 10% (rms). Values obtained by different methods also differ by around 10% (rms). Figure 10 shows an example of phase space trajectory normalization using model Twiss parameters (in black) and measured ones estimated phase data (in red). In both cases horizontal and vertical trajectories deviate from circles. To perform full normalization, coupling should be included.
@@ IX.2 Coupled Twiss Parameters
Figure 9 shows the estimated in-plane beta functions obtained using methods described in Sections VI, VII, and VIII. For comparison, beta functions estimated using the method based on TbT phases are also shown. Figure 9 (b. and d.) specifically shows the deviation of in-plane beta functions relative to the values estimated based on phases. All methods for estimating coupled Twiss parameters use reconstructed momentum, which was computed from LS fit using two monitors around the probed monitor and modeled transport matrices. In all cases, the first 128 turns were used from a single TbT measurement. All methods demonstrate similar results when compared to the method based on phases. This agreement also indicates that the Twiss parameters specific to coupling, estimated using these methods, are expected to be meaningful and can be used for coupling characterization. Estimated linear invariants based on different methods are shown in Figure 7.
@@
Figure 9: (Color) Comparison of in-plane beta function estimated using different methods based on momentum reconstruction for horizontal (a.) and vertical (c.) planes. Corresponding deviations from the beta functions obtained using phase data $\Delta=(\beta^{\varphi}-\beta)/\beta^{\varphi}$ are shown in (b.) and (d.).
Estimated uncertainties for the results obtained with the $M^{k}$ fit method are several times smaller than those for the other two methods. This is related to the optimal selection of $k$ power, as described in [18]. The optimal value of $k$ for these experimental data is 10. If computed with a non-optimal value, for example, $k=1$, the values of uncertainties rise to those of the other methods. Additionally, at the IP BPMs, the estimated Twiss parameters are significantly off, exhibiting large uncertainties. The symplectified matrix norm errors, $|\hat{M}^{k}-M^{k}|/|M^{k}|$, are around 5% (rms).
@@
Figure 10: (Color) Normalized phase space trajectory (first 256 turns are shown) at STP4 monitor computed using the model normalization matrix compared with the trajectory using the measured normalization matrix based on phase data (a. and b.), reconstruction of the one–turn matrix (c. and d.), complex normalized spectrum (e. and f.), and coupled linear invariants (g. and h.).
In general, all methods are in good agreement when comparing estimated values of Twiss parameters. For example, the covariance matrix constructed from the values of in-plane beta functions obtained by different methods has all its elements close in magnitude, which indicates a similar variation of the results obtained with different methods. Figure 7 shows invariant values computed using the estimated normalization matrix, which is based on the complex normalized spectrum. In normalized coordinates, $I=1/2\left(Q^{2}+P^{2}\right)$ is a linear invariant that can be transformed to the laboratory frame using Equation 28. For the invariant fit method, the values of invariants are directly fitted. The coupled invariants, computed by different methods, are in agreement. Also, it can be noted that these invariants differ from the uncoupled estimation.
Using the full set of coupled Twiss parameters, one can build a coupled normalization matrix and use it to normalize measured phase space trajectories. Figure 10 shows an example of such normalization using both uncoupled and coupled normalization matrices. In the case of the uncoupled normalization matrix, there is a strong deviation from a circle due to introduced coupling. In all cases, the shape of the horizontal phase space trajectory is close to a circle after coupled normalization. The latter is further evidence that the considered methods produce sensible results. For the $M^{k}$ fit method, the result for the vertical plane (Figure 10 d.) has more residual smear compared to other two methods. Overall, the normalization performance for all methods is much better than that of the model or the method based on TbT phases.
@@
Another option to check the validity of the estimated coupled Twiss parameters is to compute the coupled amplitudes $a_{x,y}$ and $a_{y,x}$ defined in Section VII. These amplitudes can be computed directly from TbT data and using estimated linear invariants and coupled Twiss parameters as follows:
@@
$a_{y,x}^{2}=2\hat{I}_{x}\left(n_{3,1}^{2}+n_{3,2}^{2}\right)$ (34)
where $n_{i,j}$ are the elements of the normalization matrix defined in Section IV. The result of this comparison is shown in Figure 11. A qualitative argument for the validity of all methods can be observed, which further indicates the correctness of the estimated coupled Twiss parameters.
Figure 11: (Color) Comparison of coupled amplitudes for horizontal (a.) and vertical (b.) planes, computed directly from TbT data (in black) and using estimated linear invariants (from fit) together with coupled Twiss parameters estimated from fit (in red), complex normalized spectrum (in blue), and reconstructed one–turn matrix (in green).
@@ X Summary
In this paper, we examine the theoretical foundations of several methods for estimating the Twiss parameters, taking into account the coupling of transverse oscillations. These methods provide complete normalization matrices at each BPM. Methods for the estimation of coupled Twiss parameters can be accessed in the reference [44]. In all methods, the reconstructed momenta, which can be computed using several BPMs and either modeled or measured transport matrices between them, are used. Since the methods directly use TbT coordinates, they are sensitive to BPM calibration errors.
All methods allow for the construction of full normalization matrices at all BPMs and do not rely on the weak coupling assumption. These matrices can be used to perform linear normalization of TbT data, compute local transport matrices, and propagate coupled Twiss parameters using model transport matrices. Estimations of linear coupled invariants are available either directly from the fit itself or via estimated normalization matrices. Statistical methods are used to estimate uncertainties due to random variations in TbT data.
The method based on reconstructing the one-turn matrix or its power uses TbT coordinates and reconstructed momenta to fit the corresponding matrix. Symplectification procedure is performed on the estimated matrix prior to the computation of coupled Twiss parameters. Another method uses normalized complex spectra and a special form of the normalization matrix to minimize an objective that depends on coupling amplitudes. This minimization procedure provides an estimation of the coupled normalization matrix.
In the last method, the objective function to be minimized involves linear coupled invariants. These invariants can be estimated from the minimization procedure, fixed by prior optimization, or computed by other means. Numerical simulations show high accuracy in determining the coupled Twiss parameters using the described methods, taking into account random BPM errors and model errors.
@@
Based on experimentally estimated Twiss parameters, it can be concluded that the methods have similar performance. The method based on the one-turn matrix fit, or its power, has smaller uncertainties when used with optimal value of power, but also exhibits more residual smear, as evident from the normalization of the phase trajectories. This method, along with the method based on the fit of coupled invariants, can combine data from several TbT measurements, for example, when the usable data length is limited. The method based on the complex normalized spectrum uses a single set of TbT data measurements. The structure of the objective function used in this method makes it less sensitive to TbT signals decoherence, since it utilizes ratios of harmonic amplitudes.
Several tests were performed to assess the validity of the estimated coupled Twiss parameters: comparison with beta functions computed based on phase data, estimation of coupled actions, normalization of phase space trajectories, and reconstruction of coupled amplitudes. The results of these tests allow us to conclude that the methods described provide an adequate description of the coupled transverse motion and can be used for its characterization.
@@ XI Acknowledgments
This work was partially supported by the Ministry of Science and Higher Education of the Russian Federation within the governmental order for Boreskov institute of Catalysis (project FWUR-2024-0041).